Write eqvations (and boUndaries ie) lsx2) / 777 toft 20 kN-m 4m 3) 1ott 260 iN Pin Conaictvon ) 10m ly000 t6ft /S1

Write shear force and bending moment equations.
Answer #1

SOLUTION

we will follow 3 simple steps in each ques

step1).... reaction calculation

step2).... sf calculation and draw SFD

step3).... BM calculation and draw BMD

1

3Mips Step L hnd reat hons 3 Kips moment about ing 3 x R- 3x RA 24 3-3x ps 332 I O 10 Sf.D a Moment Ben ding SF at at x 10 3-

2

20 k- m For Egwilibin londihon momnt at A- o Rex9 +20 - RA +S kw 2- S2 kw 1 5/4hw SE = m Rax 1-24

3

400K 400 ku wt. SPiueT Y000 Ki-t af. 4o0X10- at dreaand 400-k

4

10m m fairy moment abovt A 200x S- Re16-G Re= 2 124 2.k 2- Rpt R 200 a 00-62.5 137-5 k 2 SF. at Aa 137.5k SF af Sm 62.kM 132

5

++ RA-R 2+1000 21- Rp =+206 SF at A 250 at end B.m at Pin Just bfore o -910s t Pih =No R.m No lad